# Maritime Cybersecurity: A Comprehensive Review

### v2

@@ 3.4. Risk Assessment Method
Risk assessment techniques offer a proactive approach to identifying and minimizing potential dangers, assist in preventing accidents, reduce potential financial losses, and comply with regulatory requirements. Unlike (Huang et al., 2023), this study does not attempt to comprehensively review every risk assessment technique used in the maritime context. Rather, we concentrate on a range of typical approaches and group them into five main categories: (1) data-driven; (2) semi-quantitative; (3) qualitative; (4) quantitative; and (5) emergent methods.
@@
The goal of quantitative methods is to provide numerical estimates of risk probabilities and consequences. These methods can be further classified into two categories: (a) artificial intelligence and machine learning techniques, such as Long Short-Term Memory (LSTM) networks (Zhang et al., 2024), and (b) probabilistic and statistical approaches, such as Bayesian networks (Baksh et al., 2018). These techniques are helpful for comparing various hazards or evaluating the efficacy of risk reduction strategies, but they typically call for a significant amount of data. Accurately estimating the financial, operational, and reputational implications of cyber incidents can be difficult for those quantitative approaches (Kalogeraki et al., 2018). Qualitative methods can be broadly classified into two categories: (a) hazard identification - Hazard and Operability Analysis (HAZOP) (Zhan et al., 2009); and (b) system-based approaches - Systems Theoretic Process Analysis (STPA) (Sultana et al., 2019). The former focuses on identifying and characterizing potential hazards and risks without necessarily quantifying them. These techniques are especially helpful for complex systems where quantification may be difficult and in the early phases of risk assessment. In order to provide a more nuanced assessment than simply qualitative methods, semi-quantitative methods mix qualitative and quantitative approaches. They typically use scoring or ranking systems without the full complexity of quantitative methods. They include (a) human factor analysis, or Human Factors Analysis and Classification System (HFACS) (Chauvin et al., 2013), and (b) failure analysis, or Failure Mode, Effects, and Criticality Analysis (FMECA) (Fu et al., 2014). Large datasets are a major component of data-driven approaches, which use them to find patterns, trends, and connections pertaining to marine threats. Examples include (a) association rules-based data mining and analysis techniques (Chen et al., 2022) and (b) geographic information system (GIS)-based geospatial analysis (Wang et al., 2022). Additionally, more recent methods are beginning to emerge, including knowledge graphs (Gan et al., 2023). Every approach has its own advantages and works well for various aspects of maritime cyber.
@@
Nonetheless, the lack of an internationally recognized standard for maritime cyber risk assessment techniques could be the source of any discrepancies. Recent efforts have aimed to establish a comprehensive framework for risk analysis in maritime systems. For instance, International Association of Classification (IACS) Unified Requirement E22 (IACS, 2022b) provides a structured approach to assessing the potential impact of failures, which is crucial for prioritizing risk mitigation efforts. IACS Recommendation 171 (IACS, 2022a) offers guidance on conducting systematic risk assessments, considering factors such as probability and consequence of potential hazards. These IACS initiatives represent significant steps toward standardization, but the complexity of maritime systems and the diversity of operational contexts continue to pose challenges. Above all, there might not be enough expertise with cybersecurity and maritime experience to perform comprehensive risk assessments (Alcaide and Llave, 2020).
3.5. Regulatory Frameworks/Laws
@@
IACS Unified Requirements E26 (IACS, 2022c) and E27 (IACS, 2022d) are comprehensive cybersecurity regulatory standards specifically tailored for the maritime industry. E26 focuses on ship-level cyber resilience, while E27 addresses the cyber resilience of individual onboard systems and equipment. They cover the entire life cycle of a vessel, from design and construction to operation and maintenance, and address key areas such as network segmentation, access controls, malware protection, and incident response. They help ship owners, system integrators, and manufacturers to implement robust cybersecurity measures, reducing the risk of cyber incidents. The International Ship and Port Facility Security (ISPS) Code (Code, 2016) establishes an international framework for detecting security threats and taking preventive measures against security incidents affecting ships or port facilities used in trade. It defines roles and responsibilities for governments, local administrations, ship and port industries at the national and international level.
@@
However, these regulations do have potential drawbacks. Their complexity can be challenging for smaller operators or those with limited cybersecurity expertise to fully understand and implement. Additionally, the rapid pace of technological change in cybersecurity may challenge the ability of these standards to keep up with new threats. These regulations can become outdated quickly in the face of evolving cyber threats and create vulnerabilities in maritime cybersecurity. There has been issues in lack of standards for new technologies (Stalmokaitė et al., 2023). This shows how regulatory frameworks can lag behind technological innovations. Besides, many regulations do not provide detailed technical standards and leave much open to interpretation. Most importantly, they do not fully account for the rapid cyber threats and leave ships vulnerable to new attack vectors. For example, the ISPS Code was primarily designed to enhance maritime security against terrorism threats. Many shipping companies focused on meeting the minimum requirements of the ISPS Code. Ships compliant with ISPS might still be vulnerable to cyber attacks, as the compliance requirements did not comprehensively cover the evolving cyber threat landscape.
3.6. Human Factor
@@
Empirical research suggests that human factors account for 65.8% to 80% of maritime incidents, with human-related events and actions being identified as the key contributing reason (Shi et al., 2021). This diverse field includes a number of important elements: 1) personal cognitive aspects, such as mental workload, situation awareness, and decision-making techniques; 2) physiological elements, including exhaustion, stress, and irregularities in the circadian rhythm; 3) Interpersonal factors, which include issues with coordination, poor communication, and team dynamics; 4) the interplay between humans and machines, especially in light of the growing automation and technological complexity of maritime operations; 5) Organizational elements, such as management procedures, safety culture, and production pressure (Di Bucchianico et al., 2016)(Galieriková, 2019)(Grech et al., 2008).
@@
A paradigm change towards more nuanced considerations of human aspects in operating methods is required due to the mounting body of evidence indicating the significant role of human factors to maritime events. Task performance is directly influenced by cognitive and physiological aspects at the microlevel. For example, it has been demonstrated that fatigue, a well-known problem in the maritime industry, impairs cognitive abilities including attention, decision-making, and situational awareness, which may result in mistakes made during navigation or a delay in reacting to urgent situations (Hetherington et al., 2006). Interpersonal dynamics and the human-machine interaction significantly influence group performance at the mesolevel. Poor interface design can cause mode confusion or automation surprises, and inadequate team interactions can contribute to mistake chains that result in incidents (Chauvin, 2011). Human factors can play a significant influence in shaping corporate safety culture and policy formulation at a macro level. On the other hand, a weak safety culture may prioritize short-term operational efficiency ahead of long-term safety considerations (Håvold and Nesset, 2009). A strong safety culture can promote proactive risk management techniques and near-miss reporting. As a result, the incorporation of organizational, structural, individual, and team perspectives has highlighted the critical role that human factors play in the field of maritime cybersecurity.
4. Maritime Cyber Attacks
@@
We review existing cyber attacks, presented in both academic research and in industry, on the various maritime systems. Such academic research may focus on in-depth studies, analyzing attack patterns, investigating novel threat vectors, and exploring theoretical vulnerabilities that may not yet have been exploited in real-world scenarios. In contrast, industry reports often detail actual incidents, offering specifics about attack methodologies, targets, and outcomes that academic research might not have access to. Stakeholders in the maritime sector typically gain from a combination of theoretical knowledge and practical experience by examining both industry reports and academic research.
4.1. Real-Life Maritime Incidents (Industry)
@@
4.1.1. Multi-dimensional Attack Taxonomy
Figure 4. Multi-dimensional Maritime Cyber Attack Taxonomy
A well-defined taxonomy based on these dimensions (as illustrated in Figure 4) can provide stakeholders in the maritime industry—shipbuilders, operators, and cybersecurity experts—with important new information about the types and extent of cyberthreats that affect the industry. A comprehensive analysis of all maritime incidents is outside the purview of this article, even though various studies have given an overview of noteworthy, publicly revealed cyber attack occurrences (Meland et al., 2021; Oruc, 2020; Schwarz et al., 2021; Senarak, 2023). The following significant factors have been taken into consideration while mapping the marine incidents into our multi-dimensional attack taxonomy:
  1. (1)
@@ 4.1.2. Evaluation of Incidents by 2014 - 2023
Cybercriminals have been targeting the maritime sector, which includes ports, ships, shipping firms, and maritime authorities, more and more in recent years. The maritime industry is becoming more and more susceptible to sophisticated cyberattacks, thus recent occurrences need to be closely examined. This section therefore concentrates on cyberattacks that have happened in the last ten years, specifically from 2014 to 2023. A thorough analysis has been conducted on the twenty chosen maritime events (see Table 2 for the evaluation and Figure 5 for the timeline). Our goal is to contribute to a better understanding of the changing cyber threat landscape by exploring particular categories of contributing elements that may have been missed or insufficiently addressed in existing literature.
@@
Figure 5. The timeline of the significant maritime cyber attack incidents with color representations
@@
Business Email Compromise (BEC) in Dubai (Henry, 2015): In 2014, Nautilus Minerals and Dubai-based marine solutions company Marine Assets Corporation (MAC) fell victim to a BEC attack, an attack that tricked Nautilus paying a \$10 million deposit intended for MAC into a fraudulent bank account.
@@
Targeted Phishing in Limassol (Birkett, 2015): A Cyprus-based shipping company suffered a targeted phishing attack that resulted in a \$644,000 loss. The attacker carefully mimicked the legitimate hunker supplier’s email communications and requested the funds to be sent to a different bank account. The fraud was discovered when the legitimate fuel company contacted the shipping company for the outstanding payment.
@@
GPS Jamming in South Korea (Wood and Kim, 2016): In 2016, South Korea experienced a series of GPS jamming attacks that affected hundreds of ship, causing significant disruptions to navigation. The attacks, believed to have originated from hackers in North Korea, highlighted the vulnerability of GPS systems.
@@
Software Flaws in U.S. Ports (Cimpanu, 2016): Navis WebAccess, a web-based application essential for real-time operational data access in ports and logistics worldwide, was discovered to have a critical SQL injection vulnerability (CVE-2016-5817). Ethical hacker ”bRpsd” publicly disclosed the vulnerability by releasing proof-of-concept (PoC) exploit code. Unfortunately, this disclosure occurred without prior notification to the vendor, potentially putting users at risk from malicious actors exploiting the vulnerability.
@@
Navigation System Attack (Blake, 2017): In February 2017, hackers seized control of a German-owned container vessel’s navigation systems for 10 hours, rendering the captain unable to maneuver. The vessel’s IT system was compromised, and onboard experts had to intervene to restore control. Industry sources suggest the attack was an attempt by pirates to divert the ship for boarding and ransom.
@@
Maersk NotPetya Attack (Mathews, 2017): In 2017, the state-backed hacker group Sandworm, infamous for the NotPetya ransomware, launched a devastating cyberattack on Ukrainian businesses and organizations. It was facilitated by a widely used tax accounting software - ME Doc, and NotPetya spread rapidly, crippling computer systems not just in Ukraine, but also at global port terminals controlled by Maersk division. The attack infiltrated nearly 45,000 devices and 4,000 servers across 600 Maersk locations worldwide. This widespread disruption resulted in an estimated \$250-300 million loss in revenue for Maersk.
@@
GPS Spoofing in the Black Sea (GNSS, 2018): In 2017, alarm bells rang out across the Black Sea when over 20 vessels reported their GPS positions as being inland at an airport, reported by U.S. Maritime Administration. This wasn’t an isolated glitch – over 20 ships reporting the same false location, coupled with their positions bouncing back and forth between the airport and their true locations, strongly suggests a deliberate and large-scale GPS spoofing attack.
@@
Port of Barcelona Attack (Ilascu, 2018): In 2018, the Port of Barcelona in Spain experienced a significant cyber attack that affected its IT systems and caused disruptions to its operations. Due to system instability caused by the compromise, cargo handling efficiency between vessels and trailers significantly decreased, causing major delays in delivery.
@@
Austal Data Breach (Crozier, 2020): In mid-October 2018, Austal, Australia’s shipbuilder and defence contractor, suffered a major data breach. Attackers purchased stolen Austal login credentials on the dark web and these credentials were used to gain access to Austal’s Australian business data management systems. The attackers attempted to extort Austal by offering to return to the stolen data in exchange for a ransom payment.
@@
GNSS Spoofing in Port of Shanghai (Zeng, 2021): In 2019, several incidents of GNSS (Global Navigation Satellite System) interference were reported in Chinese coastal areas, including the Port of Shanghai. This interference can disrupt or falsify GNSS signals, potentially affecting a vessel’s navigation and communication equipment. It has been proposed that the Chinese government might be behind the GPS spoofing incidents, either as a security measure to conceal oil terminals and important government facilities or to evade surveillance of oil imports.
@@
James Fisher and Sons (JFS) (Executive, 2019): UK-based marine services provider JFS disclosed an unauthorized intrusion into its computer systems. The cyberattack triggered a 5.7% decline in the company’s market shares.
@@
AIS Spoofing in Polish Waters (Bergman, 2021): In November 2020, the Automatic Identification System (AIS) location data for the USS Roosevelt was spoofed. Threat actors, potentially linked to Russia, manipulated the AIS data to show the vessel near the Russian enclave of Kaliningrad, when it was actually elsewhere. The false AIS data could serve to create a narrative portraying Russia as the victim of encroaching naval activity.
@@
Death Kitty Ransomware (Goud, 2021): Transnet, a South Africa-based transport company, recently fell victim to a Death Kitty ransomware attack targeting its computer and NAVIS systems. This cyber incident has resulted in significant disruptions to Transnet’s operations, with potential impacts lasting up to a week. The attack specifically targeted the company’s port operations, leading to a complete seizure of port activities and a halt in the movement of cargo until the affected systems are fully restored.
@@
Attack on Greek Companies (Bader, 2021): In 2011, the companies impacted by the cyberattack leveraged the communication systems provided by Danaos Management Consultants, based in the coastal community of Piraeus in Greater Athens. This attack disrupted their ability to communicate with various stakeholders, including ships, suppliers, agents, charterers, and other key partners.
@@
Attack on the Port of Houston (Paganini, 2021): The Port of Houston in the USA encountered a cyber attack targeting its computer network. The attackers attempted to exploit a zero-day vulnerability, indicating that the flaw was not previously known to the software creator. The nature of the attack suggested the involvement of a nation-state actor, whose objective was to obtain sensitive government information and potentially disrupt or halt operations.
@@
Attack on Voyager Worldwide (Chambers, 2022): In 2022, Voyager Worldwide, a leading maritime technology company, fell victim to a cyber attack that resulted in the complete shutdown of its navigation services and solutions. This major incident affected over 1,000 shipping companies worldwide that rely on Voyager’s technology and services.
@@
Ransomware Attack on the Port of Lisbon (Ribeiro, 2023): The Port of Lisbon suffered a crippling ransomware attack and data breach on Christmas Day 2022. The LockBit group, known for their aggressive tactics, stole a trove of sensitive data including financial reports, cargo information, and ship logs. The Port of Lisbon has not publicly commented on the attack or whether they paid the ransom demands of \$1.5 million.
@@
DDoS Attack on London Port Authority Websites (Glover, 2022): The Port of London Authority (PLA) fell victim to a cyberattack which has knocked its website offline in May 2022. This incident disrupted the normal online operations and services provided by the PLA, impacting its ability to communicate important information, handle inquiries, and facilitate digital transactions.
@@
Ransomware Attack on the Port of Nagoya (Lyngaas, 2023): In July 2023, the bustling port of Nagoya in Japan faced a crippling ransomware assault orchestrated by the Lockbit group. The cyberattack effectively paralyzed the port’s computerized container handling system, halting the flow of incoming shipping containers for a duration of two days.
@@
DP World Australia Attack (Liang, 2023): In November 2023, DP World, a key player managing 40% of Australia’s maritime freight, announced a suspension of operations at its port terminals in Sydney, Melbourne, Brisbane, and Fremantle due to a cybersecurity breach. Responding to the cyber threat, the company took immediate action by isolating its computer systems from external networks, effectively shutting down operations. This resulted in a standstill, with approximately 30,000 shipping containers being left stranded.
4.1.3. Lesson Learned Summary

### v3

@@ II-D Risk Assessment Method
Risk assessment techniques offer a proactive approach to identifying and minimizing potential dangers, assist in preventing accidents, reduce potential financial losses, and comply with regulatory requirements. Unlike [35], this study does not attempt to comprehensively review every risk assessment technique used in the maritime context. Rather, we concentrate on a range of typical approaches and group them into five main categories: (1) data-driven; (2) semi-quantitative; (3) qualitative; (4) quantitative; and (5) emergent methods.
@@
The goal of quantitative methods is to provide numerical estimates of risk probabilities and consequences. These methods can be further classified into two categories: (a) artificial intelligence and machine learning techniques, such as Long Short-Term Memory (LSTM) networks [36], and (b) probabilistic and statistical approaches, such as Bayesian networks [37]. These techniques are helpful for comparing various hazards or evaluating the efficacy of risk reduction strategies, but they typically call for a significant amount of data. Accurately estimating the financial, operational, and reputational implications of cyber incidents can be difficult for those quantitative approaches [38]. Qualitative methods can be broadly classified into two categories: (a) hazard identification - Hazard and Operability Analysis (HAZOP) [39]; and (b) system-based approaches - Systems Theoretic Process Analysis (STPA) [40]. The former focuses on identifying and characterizing potential hazards and risks without necessarily quantifying them. These techniques are especially helpful for complex systems where quantification may be difficult and in the early phases of risk assessment. In order to provide a more nuanced assessment than simply qualitative methods, semi-quantitative methods mix qualitative and quantitative approaches. They typically use scoring or ranking systems without the full complexity of quantitative methods. They include (a) human factor analysis, or Human Factors Analysis and Classification System (HFACS) [41], and (b) failure analysis, or Failure Mode, Effects, and Criticality Analysis (FMECA) [42]. Large datasets are a major component of data-driven approaches, which use them to find patterns, trends, and connections pertaining to marine threats. Examples include (a) association rules-based data mining and analysis techniques [43] and (b) geographic information system (GIS)-based geospatial analysis [44]. Additionally, more recent methods are beginning to emerge, including knowledge graphs [45]. Every approach has its own advantages and works well for various aspects of maritime cyber.
@@
Nonetheless, the lack of an internationally recognized standard for maritime cyber risk assessment techniques could be the source of any discrepancies. Recent efforts have aimed to establish a comprehensive framework for risk analysis in maritime systems. For instance, International Association of Classification (IACS) Unified Requirement E22 [46] provides a structured approach to assessing the potential impact of failures, which is crucial for prioritizing risk mitigation efforts. IACS Recommendation 171 [47] offers guidance on conducting systematic risk assessments, considering factors such as probability and consequence of potential hazards. These IACS initiatives represent significant steps toward standardization, but the complexity of maritime systems and the diversity of operational contexts continue to pose challenges. Above all, there might not be enough expertise with cybersecurity and maritime experience to perform comprehensive risk assessments [48].
II-E Regulatory Frameworks/Laws
@@
IACS Unified Requirements E26 [49] and E27 [50] are comprehensive cybersecurity regulatory standards specifically tailored for the maritime industry. E26 focuses on ship-level cyber resilience, while E27 addresses the cyber resilience of individual onboard systems and equipment. They cover the entire life cycle of a vessel, from design and construction to operation and maintenance, and address key areas such as network segmentation, access controls, malware protection, and incident response. They help ship owners, system integrators, and manufacturers to implement robust cybersecurity measures, reducing the risk of cyber incidents. The International Ship and Port Facility Security (ISPS) Code [51] establishes an international framework for detecting security threats and taking preventive measures against security incidents affecting ships or port facilities used in trade. It defines roles and responsibilities for governments, local administrations, ship and port industries at the national and international level.
@@
However, these regulations do have potential drawbacks. Their complexity can be challenging for smaller operators or those with limited cybersecurity expertise to fully understand and implement. Additionally, the rapid pace of technological change in cybersecurity may challenge the ability of these standards to keep up with new threats. These regulations can become outdated quickly in the face of evolving cyber threats and create vulnerabilities in maritime cybersecurity. There has been issues in lack of standards for new technologies [52]. This shows how regulatory frameworks can lag behind technological innovations. Besides, many regulations do not provide detailed technical standards and leave much open to interpretation. Most importantly, they do not fully account for the rapid cyber threats and leave ships vulnerable to new attack vectors. For example, the ISPS Code was primarily designed to enhance maritime security against terrorism threats. Many shipping companies focused on meeting the minimum requirements of the ISPS Code. Ships compliant with ISPS might still be vulnerable to cyber attacks, as the compliance requirements did not comprehensively cover the evolving cyber threat landscape.
II-F Human Factor
@@
Empirical research suggests that human factors account for 65.8% to 80% of maritime incidents, with human-related events and actions being identified as the key contributing reason [53]. This diverse field includes a number of important elements: 1) personal cognitive aspects, such as mental workload, situation awareness, and decision-making techniques; 2) physiological elements, including exhaustion, stress, and irregularities in the circadian rhythm; 3) Interpersonal factors, which include issues with coordination, poor communication, and team dynamics; 4) the interplay between humans and machines, especially in light of the growing automation and technological complexity of maritime operations; 5) Organizational elements, such as management procedures, safety culture, and production pressure [54][55][56].
@@
Task performance is directly influenced by cognitive and physiological aspects at the microlevel. For example, it has been demonstrated that fatigue, a well-known problem in the maritime industry, impairs cognitive abilities including attention, decision-making, and situational awareness, which may result in mistakes made during navigation or a delay in reacting to urgent situations [57]. Interpersonal dynamics and the human-machine interaction significantly influence group performance at the mesolevel. Poor interface design can cause mode confusion or automation surprises, and inadequate team interactions can contribute to mistake chains that result in incidents [58]. Human factors can play a significant influence in shaping corporate safety culture and policy formulation at a macro level. On the other hand, a weak safety culture may prioritize short-term operational efficiency ahead of long-term safety considerations [59]. A strong safety culture can promote proactive risk management techniques and near-miss reporting. As a result, the incorporation of organizational, structural, individual, and team perspectives has highlighted the critical role that human factors play in the field of maritime cybersecurity.
III Maritime Cyber Attacks
@@
Academic research may focus on in-depth studies, analyzing attack patterns, investigating novel threat vectors, and exploring theoretical vulnerabilities that may not yet have been exploited in real-world scenarios. In contrast, industry reports often detail actual incidents, offering specifics about attack methodologies, targets, and outcomes that academic research might not have access to. We review existing cyber attacks, presented in both academic research and in industry, on the various maritime systems.
III-A Real-Life Maritime Incidents (Industry)
@@
III-A1 Multi-dimensional Attack Taxonomy
Figure 4: Multi-dimensional Maritime Cyber Attack Taxonomy
A well-defined taxonomy based on these dimensions (See Figure 4) can provide stakeholders in the maritime industry—shipbuilders, operators, and cybersecurity experts—with important new information about the types and extent of cyberthreats that affect the industry. A comprehensive analysis of all maritime incidents is outside the purview of this article, even though various studies have given an overview of noteworthy, publicly revealed cyber attack occurrences [60, 61, 62, 63]. The following significant factors have been taken into consideration while mapping the marine incidents into our multi-dimensional attack taxonomy:
@@ III-A2 Evaluation of Incidents by 2014 - 2023
Cybercriminals have been targeting the maritime sector, which includes ports, ships, shipping firms, and maritime authorities, more and more in recent years. The maritime industry is becoming more and more susceptible to sophisticated cyberattacks, thus recent occurrences need to be closely examined. This section therefore concentrates on cyberattacks that have happened in the last ten years, specifically from 2014 to 2023. A thorough analysis has been conducted on the twenty chosen maritime events (see Table II for the evaluation and Figure 5 for the timeline). Our goal is to contribute to a better understanding of the changing cyber threat landscape by exploring particular categories of contributing elements that may have been missed or insufficiently addressed in existing literature.
@@
Figure 5: The timeline of the significant maritime cyber attack incidents with color representations
@@
Business Email Compromise (BEC) in Dubai [64]: In 2014, Nautilus Minerals and Dubai-based marine solutions company Marine Assets Corporation (MAC) fell victim to a BEC attack, an attack that tricked Nautilus paying a \$10 million deposit intended for MAC into a fraudulent bank account.
@@
Targeted Phishing in Limassol [65]: A Cyprus-based shipping company suffered a targeted phishing attack that resulted in a \$644,000 loss. The attacker carefully mimicked the legitimate hunker supplier’s email communications and requested the funds to be sent to a different bank account. The fraud was discovered when the legitimate fuel company contacted the shipping company for the outstanding payment.
@@
GPS Jamming in South Korea [66]: In 2016, South Korea experienced a series of GPS jamming attacks that affected hundreds of ship, causing significant disruptions to navigation. The attacks, believed to have originated from hackers in North Korea, highlighted the vulnerability of GPS systems.
@@
Software Flaws in U.S. Ports [67]: Navis WebAccess, a web-based application essential for real-time operational data access in ports and logistics worldwide, was discovered to have a critical SQL injection vulnerability (CVE-2016-5817). Ethical hacker ”bRpsd” publicly disclosed the vulnerability by releasing proof-of-concept (PoC) exploit code. Unfortunately, this disclosure occurred without prior notification to the vendor, potentially putting users at risk from malicious actors exploiting the vulnerability.
@@
Navigation System Attack [68]: In February 2017, hackers seized control of a German-owned container vessel’s navigation systems for 10 hours, rendering the captain unable to maneuver. The vessel’s IT system was compromised, and onboard experts had to intervene to restore control. Industry sources suggest the attack was an attempt by pirates to divert the ship for boarding and ransom.
@@
Maersk NotPetya Attack [69]: In 2017, the state-backed hacker group Sandworm, infamous for the NotPetya ransomware, launched a devastating cyberattack on Ukrainian businesses and organizations. It was facilitated by a widely used tax accounting software - ME Doc, and NotPetya spread rapidly, crippling computer systems not just in Ukraine, but also at global port terminals controlled by Maersk division. The attack infiltrated nearly 45,000 devices and 4,000 servers across 600 Maersk locations worldwide. This widespread disruption resulted in an estimated \$250-300 million loss in revenue for Maersk.
@@
GPS Spoofing in the Black Sea [70]: In 2017, alarm bells rang out across the Black Sea when over 20 vessels reported their GPS positions as being inland at an airport, reported by U.S. Maritime Administration. This wasn’t an isolated glitch – over 20 ships reporting the same false location, coupled with their positions bouncing back and forth between the airport and their true locations, strongly suggests a deliberate and large-scale GPS spoofing attack.
@@
Port of Barcelona Attack [71]: In 2018, the Port of Barcelona in Spain experienced a significant cyber attack that affected its IT systems and caused disruptions to its operations. Due to system instability caused by the compromise, cargo handling efficiency between vessels and trailers significantly decreased, causing major delays in delivery.
@@
Austal Data Breach [72]: In mid-October 2018, Austal, Australia’s shipbuilder and defence contractor, suffered a major data breach. Attackers purchased stolen Austal login credentials on the dark web and these credentials were used to gain access to Austal’s Australian business data management systems. The attackers attempted to extort Austal by offering to return to the stolen data in exchange for a ransom payment.
@@
GNSS Spoofing in Port of Shanghai [73]: In 2019, several incidents of GNSS (Global Navigation Satellite System) interference were reported in Chinese coastal areas, including the Port of Shanghai. This interference can disrupt or falsify GNSS signals, potentially affecting a vessel’s navigation and communication equipment. It has been proposed that the Chinese government might be behind the GPS spoofing incidents, either as a security measure to conceal oil terminals and important government facilities or to evade surveillance of oil imports.
@@
James Fisher and Sons (JFS) [74]: UK-based marine services provider JFS disclosed an unauthorized intrusion into its computer systems. The cyberattack triggered a 5.7% decline in the company’s market shares.
@@
AIS Spoofing in Polish Waters [75]: In November 2020, the Automatic Identification System (AIS) location data for the USS Roosevelt was spoofed. Threat actors, potentially linked to Russia, manipulated the AIS data to show the vessel near the Russian enclave of Kaliningrad, when it was actually elsewhere. The false AIS data could serve to create a narrative portraying Russia as the victim of encroaching naval activity.
@@
Death Kitty Ransomware [76]: Transnet, a South Africa-based transport company, recently fell victim to a Death Kitty ransomware attack targeting its computer and NAVIS systems. This cyber incident has resulted in significant disruptions to Transnet’s operations, with potential impacts lasting up to a week. The attack specifically targeted the company’s port operations, leading to a complete seizure of port activities and a halt in the movement of cargo until the affected systems are fully restored.
@@
Attack on Greek Companies [77]: In 2011, the companies impacted by the cyberattack leveraged the communication systems provided by Danaos Management Consultants, based in the coastal community of Piraeus in Greater Athens. This attack disrupted their ability to communicate with various stakeholders, including ships, suppliers, agents, charterers, and other key partners.
@@
Attack on the Port of Houston [78]: The Port of Houston in the USA encountered a cyber attack targeting its computer network. The attackers attempted to exploit a zero-day vulnerability, indicating that the flaw was not previously known to the software creator. The nature of the attack suggested the involvement of a nation-state actor, whose objective was to obtain sensitive government information and potentially disrupt or halt operations.
@@
Attack on Voyager Worldwide [79]: In 2022, Voyager Worldwide, a leading maritime technology company, fell victim to a cyber attack that resulted in the complete shutdown of its navigation services and solutions. This major incident affected over 1,000 shipping companies worldwide that rely on Voyager’s technology and services.
@@
Ransomware Attack on the Port of Lisbon [80]: The Port of Lisbon suffered a crippling ransomware attack and data breach on Christmas Day 2022. The LockBit group, known for their aggressive tactics, stole a trove of sensitive data including financial reports, cargo information, and ship logs. The Port of Lisbon has not publicly commented on the attack or whether they paid the ransom demands of \$1.5 million.
@@
DDoS Attack on London Port Authority Websites [81]: The Port of London Authority (PLA) fell victim to a cyberattack which has knocked its website offline in May 2022. This incident disrupted the normal online operations and services provided by the PLA, impacting its ability to communicate important information, handle inquiries, and facilitate digital transactions.
@@
Ransomware Attack on the Port of Nagoya [82]: In July 2023, the bustling port of Nagoya in Japan faced a crippling ransomware assault orchestrated by the Lockbit group. The cyberattack effectively paralyzed the port’s computerized container handling system, halting the flow of incoming shipping containers for a duration of two days.
@@
DP World Australia Attack [83]: In November 2023, DP World, a key player managing 40% of Australia’s maritime freight, announced a suspension of operations at its port terminals in Sydney, Melbourne, Brisbane, and Fremantle due to a cybersecurity breach. Responding to the cyber threat, the company took immediate action by isolating its computer systems from external networks, effectively shutting down operations. This resulted in a standstill, with approximately 30,000 shipping containers being left stranded.
III-A3 Lesson Learned Summary